OpenStudy (anonymous):

Simplify -4 w little 6 would it be (-4)(-4)(-4)(-4)(-4)(-4)

OpenStudy (anonymous):

sorry didnt know how to write the six

OpenStudy (jlastino):

You can write an exponent in text using ^ (this means a number raised to) So -4^6 (-4 is raised to an even exponent so the resulting number would be positive) Since its positive it can also be written as 4^6 To make things easier you can use the law of exponents to express that as the product of 4w^3 and 4w^3 4^3=64 64w^3(64w^3)= ?

OpenStudy (anonymous):

what is the 3 and the w for

OpenStudy (jlastino):

w is the variable for 4 since you raised it to 6 (w^6 )

OpenStudy (jlastino):

or did you mean w/ (with)?

OpenStudy (anonymous):

yes it was with

OpenStudy (jlastino):

oh haha so it's only (64)(64)

OpenStudy (anonymous):

can I ask how we get 64

OpenStudy (jlastino):

we cubed 4 (meaning we multiply 4 by itself 3 times) 4*4*4 = 64

OpenStudy (anonymous):

so the expression x.x.x.x.x.x would be x\[^{6}\]

OpenStudy (jlastino):

Yup

OpenStudy (jlastino):

exponents were made to make writing large numbers easier haha imagine writing down x^50 lol

OpenStudy (anonymous):

yeh right lol thanks. Sorry I am new to algebra

OpenStudy (jlastino):

no prob you'll eventually get the hang of it :D